A stock issues a dividend of $1. If this dividend is expected to grow 5% per year indefinitely, and the cost of capital is 12%, what is the value of the stock?, rounded to the nearest cent?

Now imagine the stock from the previous question won't issue the dividend until three years from now. What is the value of the stock today, rounded to the nearest cent?
Answer #1

Answer a.

Last Dividend, D0 = $1.00
Growth Rate, g = 5.00%
Cost of Capital, rs = 12.00%

Expected Dividend, D1 = D0 * (1 + g)
Expected Dividend, D1 = $1.00 * 1.05
Expected Dividend, D1 = $1.05

Current Price, P0 = D1 / (rs - g)
Current Price, P0 = $1.05 / (0.12 - 0.05)
Current Price, P0 = $1.05 / 0.07
Current Price, P0 = $15.00

Answer b.

Dividend in 3 years, D3 = $1.00
Growth Rate, g = 5.00%
Cost of Capital, rs = 12.00%

Price in 2 years, P2 = D3 / (rs - g)
Price in 2 years, P2 = $1.00 / (0.12 - 0.05)
Price in 2 years, P2 = $1.00 / 0.07
Price in 2 years, P2 = $14.28571

Current Price, P0 = P2 / (1 + rs)^2
Current Price, P0 = $14.28571 / 1.12^2
Current Price, P0 = $11.39